6 numbers have average 9, if the 6th number is deducted the resulting average is 10. What is the 6th number?
Answer by mananth(16946)

6 numbers average =9
total = 6*9 =54
let the 6th number be x
(54-x)/5 = 10
54-x=50
x=54-50
x=4
6th number is 4